OpenStudy (xxlyh):

Medals! What is the value of z in the equation 2(4z – 9 – 11) = 166 – 46? 15 20 30 35

rebeccaxhawaii (rebeccaxhawaii):

any ideas?

OpenStudy (xxlyh):

no idea :/ I'm reviewing this from an old lesson I got wrong but it never gave me the correct answer

OpenStudy (xxlyh):

@rebeccaxhawaii

OpenStudy (anonymous):

combine the subtraction inside parentheses with the -9 and -11 and on the right side with 166 and -46 first

OpenStudy (anonymous):

z = 20

OpenStudy (xxlyh):

how did you get that answer?

OpenStudy (fortytherapper):

Do you know the distribution property?

OpenStudy (anonymous):

expand 2(4z-9-11). You'll get 8z-18-22

OpenStudy (anonymous):

-18-22=-40. So that is 8z-40=166-46

OpenStudy (anonymous):

166-46=120 adding 40 to both sides, we get 8z-40+40=120+40

OpenStudy (anonymous):

which is 8z=160 divide both sides by 8, we have z=20